Northeastern Insurance Company is considering opening an office in the U.S. The two cities under consideration are Philadelphia and New York. The factor ratings​ (higher scores are​ better) for the two cities are given in the following table.

     Factor   Weight   Philadelphia   New York
Customer convenience   0.25   65   75
Bank accessibility   0.20   35   95
Computer support   0.20   85   75
Rental costs   0.15   90   55
Labor costs   0.10   75   45
Taxes   0.10   90   55                                                                                                 

Factor Weight Philadelphia New York Customer convenience 0.25 65 75 Bank accessibility 0.20 35 95 Computer support 0.20 85 75 Rental costs 0.15 90 55 Labor costs 0.10 75 45 Taxes 0.10 90 55 Based on the given​ information, the best location for Northeastern Insurance Company to open the office is ▼ __________ ? New york or Phidadelphia

​, with a total weighted score of _____. ​(Enter your response rounded to two decimal​ places.)
Answer #1

Total score for any location = SUM of Weight*Factors

Philadelphia = 0.25*65 + 0.2*35 + 0.2*85 + 0.15*90 + 0.1*75 + 0.1*90 = 70.25

New York = 0.25*75 + 0.2*95 + 0.2*75 + 0.15*55 + 0.1*45 + 0.1*55 = 71.00

Ans = New York is is better due to higher score with total score of 71.00